# Science Out of Its Ivory Tower: Improving Accessibility with Reinforcement Learning

### v1

@@ 1.1 Accessible Language in Science Communication
At first glance, the daily publication of tens of thousands of scientific papers—many freely accessible through open science and open access initiatives—suggests few barriers to knowledge dissemination. However, two key facts challenge this perception and show that significant barriers remain. A recent survey by the U.S. Department of Education found that more than half of U.S. adults aged 16 to 74 (54%, or 130 million people) read at or below a sixth-grade level (Rothwell, 2020). Meanwhile, an analysis of the readability of biomedical research abstracts published from 1881 to 2015 found that scientific writing has become increasingly hard to read over time (Plavén-Sigray et al., 2017). Even when intended to be accessible, scientific abstracts typically require a postgraduate level of reading comprehension due to jargon use and sentence structure (Wang and Clark, 2024). This discrepancy leaves a significant portion of the population—including young readers and adults without advanced degrees—unable to fully engage with scientific works, even if these are made freely available online. The “infodemic” surrounding COVID-19 highlighted this issue: the urgent need for understandable information about the virus clashed with the complex presentation of scientific findings—leading many to turn to more digestible but less reliable narratives on social media (Wang et al., 2019; Islam et al., 2020; Calleja et al., 2021).
@@
While the legal and medical fields have long been encouraged to use accessible language as a clear conduit for public engagement (Mazur, 2000; Petelin, 2010), momentum for the adoption of accessible language within scientific communities has been building, roughly since the start of the open science movement (Schriver, 2017). For instance, the National Institutes of Health (NIH) advocates for “clear and simple” principles when communicating with audiences with limited health literacy, and the Proceedings of the National Academy of Sciences of the United States of America (PNAS) requires authors to submit a significance statement accessible to non-experts (Berenbaum, 2021; Pool et al., 2021). However, there are inherent conflicts between the specialized nature of communication among disciplinary peer scholars and the public-oriented dissemination of scientific findings. Even assuming that communicating scholarly works in plain language is possible, it will inevitably increase the communication cost among domain experts and create confusion at the more advanced levels, compared to the use of jargon and technical terms. In an era of increasingly specialized scientific research, this conundrum is not easily addressed by scientists or disseminators. Hence, given the current landscape, the widespread adoption of accessible language in scholarly works is unlikely in the near future.
@@
In response, we propose addressing the need for communicating scientific findings to a broader audience by rewriting scholarly abstracts with simpler words and grammar using language models. Since readability is key to comprehending scholarship (Flesch, 1946; DuBay, 2004; Kerwer et al., 2021), we envision the resulting accessible narratives as paving the way for the “last mile” of science, broadening access to scientific understanding and engagement, especially for younger readers and those without a college degree.
@@ 1.2 Challenges to Effective Simplification
Fine-tuning a language model using pairs of abstracts and their accessible versions is the de facto method for automating the rewriting of scholarly abstracts into more accessible versions (Xu et al., 2015; Goldsack et al., 2022; Joseph et al., 2023). Accordingly, we introduced the Scientific Abstract-Significance Statement (SASS) corpus (Wang and Clark, 2024), a dataset composed of paired abstracts and significance statements from diverse disciplines, with the latter targeting “an undergraduate-educated scientist outside their field of specialty” (Berenbaum, 2021; Pool et al., 2021). Although the simplified abstracts generated from language models fine-tuned on the SASS corpus are approximately three grade levels more readable than the original abstracts, as measured by U.S. grade-based readability scores (Wang and Clark, 2024, Sec. 6), the documents are still not sufficiently accessible; even the best models produce college-level texts. Additionally, because the vocabulary used in significance statements is often just as complex as that found in the abstracts themselves (Wang and Clark, 2024, Sec. 3), the readability improvements are primarily due to shorter sentences, and technical terms remain inadequately addressed.
Alternatively, the optimization of a language model can be guided by a chosen objective in an actor-critic manner (Ramamurthy et al., 2023). It is intuitive to choose an established document readability measure, such as the Automated Readability Index (ARI; see Section 5.2.2), to assess the overall readability of the outputs generated by the language model.11endnote: 1The objective can also be reading difficulties ranked by human raters, similar to reinforcement learning from human feedback (RLHF) (Ouyang et al., 2022), where the feedback focuses specifically on document readability rather than general preferences. While this type of feedback is of high quality, it is significantly more expensive than our reward functions. Additionally, document readability is more tangible and therefore easier to quantify than the more ambiguous concept of “human preference.” For these reasons, we did not consider using human raters in this study. However, we found that the optimization of language models guided by ARI is highly unstable, often resulting in the production of seemingly more accessible versions that still contain many technical terms. Inspired by Riddell and Igarashi (2021), we decomposed the measurement of document readability into two distinct measures: one at the sentence level and one at the word level. We then prioritized word-level accessibility in the optimization to encourage the model to use more accessible words instead of simply shortening sentences.
@@ 2.1 The Common Challenges of Science Communication
Science communication research uses several models to describe strategies for public understanding of and engagement with science, including the deficit model, the dialogue model, and the participatory model (Trench, 2008; Hetland, 2014; Sokolovska et al., 2019). The deficit model assumes that the public lacks scientific knowledge and that science communication should focus on simplifying complex scholarship through one-way information transmission. The dialogue model encourages two-way communication, where scientists and the public engage in conversations that foster mutual understanding (Trench, 2008; Joly and Kaufmann, 2008). The participatory model involves the public as active participants in the scientific process, recognizing their contributions as equally valuable to those of scientists (Brossard and Lewenstein, 2009). In practice, the dialogue model suggests that researchers should engage with the public through mediums such as social media (Davies, 2008; Hara et al., 2019; Knox and Hara, 2021), by crafting more digestible manuscripts in research (Maurer et al., 2021), and by developing plain language guidelines for practitioners (Grene et al., 2017). Communicators are encouraged to scaffold content by providing the audience with relevant, easily understandable material that either gradually increases in complexity or includes simplified explanations, examples, and references to better support the learning process (Wolfe, 2008; Landrum et al., 2014). These models often coexist and overlap in practice (Brossard and Lewenstein, 2009; Metcalfe, 2022), with practitioners employing multiple strategies to increase the readability of complex information.
In regular practice, two notable challenges persist for researchers when attempting to effectively communicate their findings to lay audiences. First, placing the burden of effective communication on scholars, who are trained in conducting research and not in effective communication, can be overwhelming. Second, translating scientific findings that heavily utilize specialized jargon often involves sacrificing some specificity in favor of accessibility. Balancing the precision required by the scientific community with the clarity needed for public understanding remains a significant challenge. However, in order to uphold the fundamental principles of open science and the goal of science communication, the public’s understanding of scholarship should be considered a requisite part of the scholarly process (Burns et al., 2003; Fecher and Friesike, 2014).
@@ 2.2 Text Simplification
The natural language processing community addresses document accessibility through automated text simplification, a process that rewrites complex documents using simpler grammar and vocabulary while retaining the original meaning (Chandrasekar et al., 1996; Alva-Manchego et al., 2020; Al-Thanyyan and Azmi, 2021). In the early days of text simplification, knowledge-based approaches were widely used (De Belder and Moens, 2010; Zhao et al., 2018; Hijazi et al., 2022). These methods relied on predefined sets of linguistic rules to transform complex sentences into simpler forms. For example, such systems would apply syntactic transformations, such as breaking down compound sentences into simpler ones, and lexical substitutions, where difficult words were replaced with more frequent, simpler synonyms. While effective in specific cases, rule-based systems were limited by the rigidity of their rules, often struggling with the variability of natural language.
@@
With the rise of machine learning, approaches that learn simplification patterns from large corpora of text have become prominent. The most common method currently is fine-tuning a pre-trained language model with parallel corpora containing paired documents of different readability levels (Xu et al., 2015; Goldsack et al., 2022; Joseph et al., 2023). These parallel corpora allow the model to “translate” complex texts into simpler ones, typically using cross-entropy loss to align the predicted simplified version with the target text. This shift to data-driven methods allows for greater flexibility and more nuanced simplifications, as the models learn from a broader range of linguistic patterns.
@@
The effectiveness that supervised fine-tuning can achieve depends on the quality of the parallel corpus, specifically whether the simplified documents are sufficiently accessible. As an example, Devaraj et al. (2021) achieved approximately a one-grade reading level improvement, comparable to the readability of the plain-language samples, by fine-tuning a language model (i.e., BART (Lewis et al., 2019)) with a corpus of pairs of technical and plain-language medical texts. This is expected: for sequence-to-sequence models, such as BART and T5, the objective of supervised fine-tuning is to minimize the Kullback–Leibler (KL) divergence between the predicted token distribution and the target distribution. Similarly, for causal language models, such as GPT and Gemma, minimizing KL divergence is equivalent to maximizing the joint probability of generating accessible sequences, conditioned on the abstracts. Consequently, the best outcome that models can achieve through supervised fine-tuning is to yield lay versions that as simple as—but no simpler than—the simplified examples in the corpus.
@@
The problems is that, in practice, high-quality corpora for simplification are either unavailable or prohibitively expensive to obtain, and widely adopted text simplification targets are often only marginally more readable than their sources. For example, the Simple English Wikipedia is, on average, less than two grade levels more readable than standard English Wikipedia articles (Isaksson, 2018). This issue also extends to abstract simplification tasks, where the simplified documents are typically less than one grade level more readable than the abstracts themselves (Wang and Clark, 2024). Carlson et al. (2018) proposed evaluating the readability of simplified texts using different versions of the Bible, which cover a wide spectrum of readability levels. This approach is specifically designed for evaluation purposes rather than for training a general-purpose simplification model.
@@
Efforts to overcome this data bottleneck have focused on two main approaches. First, scholars have modified the cross-entropy objective to achieve additional improvements, often by incorporating knowledge on word-level difficulty (Nishihara et al., 2019; Devaraj et al., 2021; Yanamoto et al., 2022). For instance, by adding pre-determined terms to the cross-entropy objective to reduce the model’s probability assigned to a set of technical words—identified using a classifier trained on essays of different readability levels—a BART model can achieve an additional one-grade reading level improvement over one trained with the standard cross-entropy objective (Devaraj et al., 2021). However, while including extra terms in the objective function can be useful, this approach tends to be rather rigid and may compromise language quality. Moreover, the performance gains from these improvements only improve readability by around one grade level. Additional performance gains have also been observed when fine-tuning larger language models (ranging from 2 to 7 billion parameters) (Wang and Clark, 2024). Having been exposed to a significantly larger number of documents, the model may land on a parameter landscape slightly better than the one represented by the target documents. Although using a larger language model results in about a three-grade-level boost in readability, it still falls approximately three grade levels short of making the abstracts generally accessible to our target audience—individuals without a college degree.
@@
Our means of addressing the data bottleneck is to optimize language models by incorporating dynamically adjusted terms into the cross-entropy objective in an actor-critic manner. This method aligns with reinforcement learning techniques that integrate multiple rewards from different types of feedback; in long-form question-answering tasks, Wu et al. (2023) have demonstrated that incorporating various reward types—such as coherence, relevance, and completeness—outperforms approaches relying on a single, holistic reward. We adapted this approach by assessing document readability through a careful balance of two heuristics proposed by Riddell and Igarashi (2021); these estimate syntactic and lexical difficulty, respectively.
@@ 3 Scientific Abstract-Significance Statement (SASS) Corpus
We used the Scientific Abstract-Significance Statement (SASS) corpus in our experiments. This corpus is composed of 3,430 abstract-significance statement pairs derived from PNAS and divided into training (3,030 samples), validation (200 samples), and test sets (200 samples) (Wang and Clark, 2024). It covers a wide range of disciplines, ensuring diverse representation across various fields, as shown in Figure 1. Corpus statistics are shown in Table 1; refer to Section 5.2 for a detailed description of the measures.
@@
Figure 1: Discipline and readability distributions of abstracts and significance statements found in the training set of the Scientific Abstract-Significance Statement corpus. The count of paired samples in different disciplines is shown in blue bars on a log10 scale (disciplines with fewer than three samples are not shown). Readability is measured using the Automated Readability Index (ARI), which estimates the number of years of schooling required to understand a text. On average, abstracts have a readability slightly below 20 ARI, indicating a post-graduate level. Significance statements are generally more readable than their corresponding abstracts.

### v2

@@ 1 Introduction
Figure 1: RLAM Training Workflow: RLAM rewrites scholarly abstracts using PPO (Section 4), guided by a reward function that balances average sentence length and word accessibility. The policy model is optimized iteratively through an actor-critic framework: it generates simplified abstracts, whose quality is assessed by the reward function and regularized through KL divergence from the frozen reference model (the supervised fine-tuning model). The reward signal is further contrasted with expected returns estimated by the value model, implemented as a linear head atop the policy model. The resulting advantage is distributed across tokens via Generalized Advantage Estimation (GAE). RLAM enables the lightweight Gemma-2B model to reduce abstracts from postgraduate to high school readability, achieving a 90% improvement over the supervised baseline.
@@ 1.1 Accessible Language in Science Communication
At first glance, the daily publication of tens of thousands of scientific papers—many freely accessible through open science and open access initiatives—suggests few barriers to knowledge dissemination. However, two key facts challenge this perception and show that significant barriers remain. A recent survey by the U.S. Department of Education found that more than half of U.S. adults aged 16 to 74 (54%, or 130 million people) read at or below a sixth-grade level (Rothwell,, 2020). Meanwhile, an analysis of the readability of biomedical research abstracts published from 1881 to 2015 found that scientific writing has become increasingly hard to read over time (Plavén-Sigray et al.,, 2017). Even when intended to be accessible, scientific abstracts typically require a postgraduate level of reading comprehension due to jargon use and sentence structure (Wang and Clark,, 2025). This discrepancy leaves a significant portion of the population—including young readers and adults without advanced degrees—unable to fully engage with scientific works, even if these are made freely available online. The “infodemic” surrounding COVID-19 highlighted this issue: the urgent need for understandable information about the virus clashed with the complex presentation of scientific findings—leading many to turn to more digestible but less reliable narratives on social media (Wang et al.,, 2019; Islam et al.,, 2020; Calleja et al.,, 2021).
@@
While the legal and medical fields have long been encouraged to use accessible language as a clear conduit for public engagement (Mazur,, 2000; Petelin,, 2010), momentum for the adoption of accessible language within scientific communities has been building, roughly since the start of the open science movement (Schriver,, 2017). For instance, the National Institutes of Health (NIH) advocates for “clear and simple” principles when communicating with audiences with limited health literacy, and the Proceedings of the National Academy of Sciences of the United States of America (PNAS) requires authors to submit a significance statement accessible to non-experts (Berenbaum,, 2021; Pool et al.,, 2021). However, there are inherent conflicts between the specialized nature of communication among disciplinary peer scholars and the public-oriented dissemination of scientific findings. Even assuming that communicating scholarly works in plain language is possible, it will inevitably increase the communication cost among domain experts and create confusion at the more advanced levels, compared to the use of jargon and technical terms. In an era of increasingly specialized scientific research, this conundrum is not easily addressed by scientists or disseminators. Moreover, expecting individual researchers to translate their findings for the public imposes an unrealistic burden, given their primary focus on advancing specialized knowledge. At the same time, public demand for clarity remains high, especially during fast-moving scientific crises where accessible information is urgently needed.
@@
In response, we propose addressing the need for communicating scientific findings to a broader audience by rewriting scholarly abstracts with simpler words and grammar using language models. Since readability is key to comprehending scholarship (Flesch,, 1946; DuBay,, 2004; Kerwer et al.,, 2021), we envision the resulting accessible narratives as paving the way for the “last mile” of science, broadening access to scientific understanding and engagement, especially for younger readers and those without a college degree.
@@ 1.2 Challenges to Effective Simplification
Fine-tuning a language model using pairs of abstracts and their accessible versions is the de facto method for automating the rewriting of scholarly abstracts into more accessible versions (Xu et al.,, 2015; Goldsack et al.,, 2022; Joseph et al.,, 2023). Accordingly, we introduced the Scientific Abstract-Significance Statement (SASS) corpus (Wang and Clark,, 2025), a dataset composed of paired abstracts and significance statements from diverse disciplines, with the latter targeting “an undergraduate-educated scientist outside their field of specialty” (Berenbaum,, 2021; Pool et al.,, 2021). Although the simplified abstracts generated from language models fine-tuned on the SASS corpus are approximately three grade levels more readable than the original abstracts, as measured by U.S. grade-based readability scores (Wang and Clark,, 2025, Sec. 6), the documents are still not sufficiently accessible; even the best models produce college-level texts. Additionally, because the vocabulary used in significance statements is often just as complex as that found in the abstracts themselves (Wang and Clark,, 2025, Sec. 3), the readability improvements are primarily due to shorter sentences, and technical terms remain inadequately addressed.
Alternatively, the optimization of a language model can be guided by a chosen objective in an actor-critic manner (Ramamurthy et al.,, 2023). It is intuitive to choose an established document readability measure, such as the Automated Readability Index (ARI; see Section 5.2.2), to assess the overall readability of the outputs generated by the language model. [^1] The objective can also be reading difficulties ranked by human raters, similar to reinforcement learning from human feedback (RLHF) (Ouyang et al.,, 2022), where the feedback focuses specifically on document readability rather than general preferences. While this type of feedback is of high quality, it is significantly more expensive than our reward functions. Additionally, document readability is more tangible and therefore easier to quantify than the more ambiguous concept of “human preference.” For these reasons, we did not consider using human raters in this study. However, we found that the optimization of language models guided by ARI is highly unstable, often resulting in the production of seemingly more accessible versions that still contain many technical terms. Inspired by Riddell and Igarashi, (2021), we decomposed the measurement of document readability into two distinct measures: one at the sentence level and one at the word level. We then prioritized word-level accessibility in the optimization to encourage the model to use more accessible words instead of simply shortening sentences.
@@ 2.1 The Common Challenges of Science Communication
Science communication research uses several models to describe strategies for public understanding of and engagement with science, including the deficit model, the dialogue model, and the participatory model (Trench,, 2008; Hetland,, 2014; Sokolovska et al.,, 2019). The deficit model assumes that the public lacks scientific knowledge and that science communication should focus on simplifying complex scholarship through one-way information transmission. The dialogue model encourages two-way communication, where scientists and the public engage in conversations that foster mutual understanding (Trench,, 2008; Joly and Kaufmann,, 2008). The participatory model involves the public as active participants in the scientific process, recognizing their contributions as equally valuable to those of scientists (Brossard and Lewenstein,, 2009). In practice, the dialogue model suggests that researchers should engage with the public through mediums such as social media (Davies,, 2008; Hara et al.,, 2019; Knox and Hara,, 2021), by crafting more digestible manuscripts in research (Maurer et al.,, 2021), and by developing plain language guidelines for practitioners (Grene et al.,, 2017). Communicators are encouraged to scaffold content by providing the audience with relevant, easily understandable material that either gradually increases in complexity or includes simplified explanations, examples, and references to better support the learning process (Wolfe,, 2008; Landrum et al.,, 2014). These models often coexist and overlap in practice (Brossard and Lewenstein,, 2009; Metcalfe,, 2022), with practitioners employing multiple strategies to increase the readability of complex information.
In regular practice, two notable challenges persist for researchers when attempting to effectively communicate their findings to lay audiences. First, placing the burden of effective communication on scholars, who are trained in conducting research and not in effective communication, can be overwhelming. Second, translating scientific findings that heavily utilize specialized jargon often involves sacrificing some specificity in favor of accessibility. Balancing the precision required by the scientific community with the clarity needed for public understanding remains a significant challenge. However, in order to uphold the fundamental principles of open science and the goal of science communication, the public’s understanding of scholarship should be considered a requisite part of the scholarly process (Burns et al.,, 2003; Fecher and Friesike,, 2014).
@@ 2.2 Text Simplification
The natural language processing community addresses document accessibility through automated text simplification, a process that rewrites complex documents using simpler grammar and vocabulary while retaining the original meaning (Chandrasekar et al.,, 1996; Alva-Manchego et al.,, 2020; Al-Thanyyan and Azmi,, 2021). In the early days of text simplification, knowledge-based approaches were widely used (De Belder and Moens,, 2010; Zhao et al.,, 2018; Hijazi et al.,, 2022). These methods relied on predefined sets of linguistic rules to transform complex sentences into simpler forms. For example, such systems would apply syntactic transformations, such as breaking down compound sentences into simpler ones, and lexical substitutions, where difficult words were replaced with more frequent, simpler synonyms. While effective in specific cases, rule-based systems were limited by the rigidity of their rules, often struggling with the variability of natural language.
@@
With the rise of machine learning, approaches that learn simplification patterns from large corpora of text have become prominent. The most common method currently is fine-tuning a pre-trained language model with parallel corpora containing paired documents of different readability levels (Xu et al.,, 2015; Goldsack et al.,, 2022; Joseph et al.,, 2023). These parallel corpora allow the model to “translate” complex texts into simpler ones, typically using cross-entropy loss to align the predicted simplified version with the target text. This shift to data-driven methods allows for greater flexibility and more nuanced simplifications, as the models learn from a broader range of linguistic patterns.
@@
The effectiveness that supervised fine-tuning can achieve depends on the quality of the parallel corpus, specifically whether the simplified documents are sufficiently accessible. As an example, Devaraj et al., (2021) achieved approximately a one-grade reading level improvement, comparable to the readability of the plain-language samples, by fine-tuning a language model (i.e., BART (Lewis et al.,, 2019)) with a corpus of pairs of technical and plain-language medical texts. This is expected: for sequence-to-sequence models, such as BART and T5, the objective of supervised fine-tuning is to minimize the Kullback–Leibler (KL) divergence between the predicted token distribution and the target distribution. Similarly, for causal language models, such as GPT and Gemma, minimizing KL divergence is equivalent to maximizing the joint probability of generating accessible sequences, conditioned on the abstracts. Consequently, the best outcome that models can achieve through supervised fine-tuning is to yield lay versions that as simple as—but no simpler than—the simplified examples in the corpus.
@@
The problems is that, in practice, high-quality corpora for simplification are either unavailable or prohibitively expensive to obtain, and widely adopted text simplification targets are often only marginally more readable than their sources. For example, the Simple English Wikipedia is, on average, less than two grade levels more readable than standard English Wikipedia articles (Isaksson,, 2018). This issue also extends to abstract simplification tasks, where the simplified documents are typically less than one grade level more readable than the abstracts themselves (Wang and Clark,, 2025). Carlson et al., (2018) proposed evaluating the readability of simplified texts using different versions of the Bible, which cover a wide spectrum of readability levels. This approach is specifically designed for evaluation purposes rather than for training a general-purpose simplification model.
@@
Efforts to overcome this data bottleneck have focused on two main approaches. First, scholars have modified the cross-entropy objective to achieve additional improvements, often by incorporating knowledge on word-level difficulty (Nishihara et al.,, 2019; Devaraj et al.,, 2021; Yanamoto et al.,, 2022). For instance, by adding pre-determined terms to the cross-entropy objective to reduce the model’s probability assigned to a set of technical words—identified using a classifier trained on essays of different readability levels—a BART model can achieve an additional one-grade reading level improvement over one trained with the standard cross-entropy objective (Devaraj et al.,, 2021). However, while including extra terms in the objective function can be useful, this approach tends to be rather rigid and may compromise language quality. Moreover, the performance gains from these improvements only improve readability by around one grade level. Additional performance gains have also been observed when fine-tuning larger language models (ranging from 2 to 7 billion parameters) (Wang and Clark,, 2025). Having been exposed to a significantly larger number of documents, the model may land on a parameter landscape slightly better than the one represented by the target documents. Although using a larger language model results in about a three-grade-level boost in readability, it still falls approximately three grade levels short of making the abstracts generally accessible to our target audience—individuals without a college degree.
@@
Our means of addressing the data bottleneck is to optimize language models by incorporating dynamically adjusted terms into the cross-entropy objective in an actor-critic manner. This method aligns with reinforcement learning techniques that integrate multiple rewards from different types of feedback; in long-form question-answering tasks, Wu et al., (2023) have demonstrated that incorporating various reward types—such as coherence, relevance, and completeness—outperforms approaches relying on a single, holistic reward. We adapted this approach by assessing document readability through a careful balance of two heuristics proposed by Riddell and Igarashi, (2021); these estimate syntactic and lexical difficulty, respectively.
@@ 3 Scientific Abstract-Significance Statement (SASS) Corpus
We used the Scientific Abstract-Significance Statement (SASS) corpus in our experiments. This corpus is composed of 3,430 abstract-significance statement pairs derived from PNAS and divided into training (3,030 samples), validation (200 samples), and test sets (200 samples) (Wang and Clark,, 2025). It covers a wide range of disciplines, ensuring diverse representation across various fields, as shown in Figure 2. Corpus statistics are shown in Table 1; refer to Section 5.2 for a detailed description of the measures.
@@
Figure 2: Discipline and readability distributions of abstracts and significance statements found in the training set of the Scientific Abstract-Significance Statement corpus. The count of paired samples in different disciplines is shown in blue bars on a log10 scale (disciplines with fewer than ten samples are not shown). Readability is measured using the Automated Readability Index (ARI), which estimates the number of years of schooling required to understand a text. On average, abstracts have a readability slightly below 20 ARI, indicating a post-graduate level. Significance statements are generally more readable than their corresponding abstracts. Orange arrows indicate the change in readability from abstracts to significance statements.